The rate of defectiveness of items from a production line is 2% five items are randomly selected from the production line
a) present the probability distribution table .
b) At least two items are defective
c) At least two items are defective
d) Given that 100 items we're randomly selected instead, find the average number of defective items
1
Expert's answer
2020-05-13T20:24:15-0400

n=5,p=0.02,q=1-0.02=0.98

a)

x-amount of defective items

P(x=0)=C5,0*0.020*0.985=0.9039

P(x=1)=C5,1*0.021*0.984=0.0922

P(x=2)=C5,2*0.022*0.983=0.0037

P(x=3)=C5,3*0.023*0.982=7.683*10-5

P(x=4)=C5,4*0.024*0.98=7.84*10-7

P(x=5)=0.025=3.2*10-9

Table

Defective 0 1 2 3 4 5

P 0.9039 0.0922 0.0037 7.683*10-5 7.84*10-7 3.2*10-9

b) At least two items are defective  

P(x≥2)=1-P(x=0)-P(x=1)=1-0.9039-0.0922=0.0039

c) At least two items are defective  

P(x≥2)=1-P(x=0)-P(x=1)=1-0.9039-0.0922=0.0039

d) μ=np=100*0.02=2